What is the slope of a line perpendicular to the line whose equation is 6x-5y=-406x−5y=−40
Answer by math_helper(2461)


6x - 5y = -40


For the given line, you can re-arrange to y=mx+b ... y=(6/5)x + 8 ... or recognize the slope of Ax+By = C is -A/B = -6/-5 = 6/5.


For a line perpendicular to the given line, the slope will be -1/(6/5) = -5/6